# MGAug: Multimodal Geometric Augmentation in Latent Spaces of Image Deformations

### v2

@@ Abstract
Geometric transformations have been widely used to augment the size of training images. Existing methods often assume a unimodal distribution of the underlying transformations between images, which limits their power when data with multimodal distributions occur. In this paper, we propose a novel model, Multimodal Geometric Augmentation (MGAug), that for the first time generates augmenting transformations in a multimodal latent space of geometric deformations. To achieve this, we first develop a deep network that embeds the learning of latent geometric spaces of diffeomorphic transformations (a.k.a. diffeomorphisms) in a variational autoencoder (VAE). A mixture of multivariate Gaussians is formulated in the tangent space of diffeomorphisms and serves as a prior to approximate the hidden distribution of image transformations. We then augment the original training dataset by deforming images using randomly sampled transformations from the learned multimodal latent space of VAE. To validate the efficiency of our model, we jointly learn the augmentation strategy with two distinct domain-specific tasks: multi-class classification on 2D synthetic dataset and segmentation on real 3D brain magnetic resonance images (MRIs). We also compare MGAug with state-of-the-art transformation-based image augmentation algorithms. Experimental results show that our proposed approach outperforms all baselines by significantly improved prediction accuracy. Our code will be publicly available on GitHub.
@@ I Introduction
Recent advances in deep learning (DL) have led to remarkable progress in image analysis tasks, including but not limited to image classification [43, 38], segmentation [9, 47, 65], object detection [31, 49], and annotation [58, 10, 67]. The literature has shown that deep neural networks (DNNs) are most powerful when trained on vast amounts of manually labeled data. Unfortunately, many image applications have relatively small annotated datasets available, due to privacy concerns, regulations, costs, and incidents of disease. The DNNs, in such cases, tend to overfit due to their poor generalizations to large data variances. Data augmentation (DA) is a simple, yet efficient and commonly used strategy to overcome this issue by synthesizing more training examples to artificially increase the dataset [55, 11, 65, 68, 53]. It is often employed as a preprocessing step applied to the training data to expose the DNNs to various types of approximated invariance; hence reducing the problem of overfitting. In the image domain, common augmentation strategies include adding variations to the image intensities [23, 18], or deforming the image by randomly generated geometric transformations (i.e., translation, rotation, or scaling) [42]. This paper focuses on DA, particularly in geometric spaces.
@@
In practice, DA is usually a manual process where a human specifies a set of affine or deformable transformations believed to make DL models invariant [22, 61]. However, such methods can pose a high risk of producing unrealistic or implausible variations of the image. For example, rotating an image 180 degrees makes sense for natural images, but produces ambiguities on digital numbers (i.e., the digits ‘6’ and ‘9’). In medical domain, randomly generated deformations may transform images from a healthy to a diseased group. To alleviate this issue, researchers have developed a group of automatic DA methods that learn an optimal augmenting policy from the data itself [11, 36, 60, 68, 2, 45]. One approach is to learn a generator via generative adversarial networks (GANs) to synthesize data from scratch [40, 1, 37]. However, these models are difficult to train due to the high complexity of largely unknown parameter spaces [28]. Despite showing encouraging results on relatively easy datasets (e.g., $28^{2}$ MNIST), they do not appear to be applicable to more complex datasets with large structural and geometric variances. Another common approach is to learn a probabilistic distribution of a set of deformable transformations derived from the training images [20, 33, 13, 65, 48, 53]. Most importantly, these methods assume a unimodal distribution of the underlying transformations between images. They do not provide an appropriate way to learn the full space of geometric transformations among the training data. This greatly limits their power when multimodal data (e.g., time-series/longitudinal images, or images of multiple sub-populations) occurs.
@@
In this paper, we present a novel model, MGAug, that for the first time learns augmenting transformations in a multimodal latent space of geometric deformations. In contrast to existing approaches limited to data with unimodal distributions, our MGAug features main advantages of
  1. 1.
@@
We demonstrate the effectiveness of our proposed model in two distinct domain-specific tasks: multi-class classification and segmentation on both 2D synthetic data and real 3D brain magnetic resonance images (MRIs), respectively. We also compare MGAug with the state-of-the-art transformation-based augmentation algorithms [48, 13]. Experimental results show that our method outperforms all baseline algorithms by significantly improved prediction accuracy.
@@ II Background: Geometric Deformations via Template-based Image Registration
$\text{with}\quad\frac{d\phi_{n}(t)}{dt}=-D\phi_{n}(t)\cdot v_{n}(t),$ (2)
@@ III Our Method: MGAug
This section presents a novel algorithm, MGAug, that learns a multimodal latent distribution of geometric transformations represented by the initial velocity fields $v_{0}$. We first introduce a generative model in latent geometric spaces for DA based on VAE, followed by an approximated variational inference. The augmented data is then used to train various image analysis tasks, i.e., image classification or segmentation. Similar to [65, 6, 7], we employ a joint learning scheme to optimize the MGAug and IAT network parameters. To simplify the notation, we omit the time index of $v_{0}$ in the following sections.
@@ III-A Generative Model in Latent Multimodal Geometric Spaces
Given a group of $N$ input images $\mathbf{J}$ with a template $I$, we assume the latent geometric transformations $\mathbf{x}$ are generated from a number of $C$ clusters. The distribution of $\mathbf{x}$ can be formulated as a Gaussian mixture with means $\boldsymbol{\mu}$ and variances $\mathbf{\Sigma}$, specified by a VAE encoder parameterized by $\beta$, i.e., $p_{\theta}(\mathbf{x}|{\boldsymbol{\epsilon}})$, where $\boldsymbol{\epsilon}$ is a random number generated from $\mathcal{N}(0,\mathbb{I})$. That is to say, the VAE encoder outputs a set of $C$ means $\boldsymbol{\mu}$ and $C$ variances $\mathbf{\Sigma}$. A one-hot vector $\mathbf{z}$ is sampled from the mixing probability $\mathbf{\boldsymbol{\pi}}\in[0,1]$ with $\sum_{c=1}^{C}\mathbf{\boldsymbol{\pi}}^{c}=1$, which chooses one component from the Gaussian mixture. In this paper, we set the parameter $\mathbf{\boldsymbol{\pi}}=C^{-1}$ to make the variable $\mathbf{z}$ uniformly distributed, which is similar to [16]. The initial velocity field $\mathbf{v}$ is generated from a VAE decoder parameterized by $\theta$ and the continuous latent variable $\mathbf{x}$. We parameterize the $\mathbf{v}$ in low-dimensional Fourier space. The entire process to generate the observed image data $\mathbf{J}$ from a set of latent variables $\{\mathbf{v},\mathbf{x},\mathbf{\boldsymbol{\epsilon}},\mathbf{z}\}$ is:
@@
Similar to [62, 63], we define the covariance matrix $\boldsymbol{\Sigma}$ as an inverse operator of the differential operator, $L$, which is used in the registration regularity in Eq. (2).
@@ III-B Variational Lower Bound
We will train the generative model in Eq. (12) with the variational inference objective, i.e. the log-evidence lower bound (ELBO). With an assumption that the mean-field variational family $q(\mathbf{v},\mathbf{x},\mathbf{\boldsymbol{\epsilon}},\mathbf{z}\,|\,\mathbf{J})$ is an approximate to the posterior, we are able to formulate the ELBO as
@@
We factorize the mean-field variational family as $\mathbf{q}=\prod_{i=1}^{n}q(\mathbf{v}_{i}|\mathbf{J}_{i},I,\boldsymbol{\lambda})\,q_{\psi_{\mathbf{x}}}(\mathbf{x}_{i}|\mathbf{v}_{i},\mathbf{J}_{i})\,q_{\psi_{\mathbf{\boldsymbol{\epsilon}}}}(\mathbf{\boldsymbol{\epsilon}}_{i}|\mathbf{v}_{i})\,p_{\beta}(\mathbf{z}_{i}|\mathbf{x}_{i},\mathbf{\boldsymbol{\epsilon}}_{i}),$ where $i$ denotes the $i$-th image. Each variational factor is parameterized with the recognition networks $\psi_{\mathbf{x}}$ and $\psi_{\mathbf{\boldsymbol{\epsilon}}}$ that output the parameters of the variational Gaussian distributions. Note that we will drop the index $i$ in the following sections for the purpose of notation simplicity.
@@
$-\mathbf{E}_{q(\mathbf{x}|\mathbf{J})\,q(\boldsymbol{\epsilon}|\mathbf{J})}[KL(p_{\beta}(\mathbf{z}|\mathbf{x},\boldsymbol{\epsilon})\,||\,p(\mathbf{z}))]$ (14)
III-C Joint Learning of MAGug With Image Analysis
We demonstrate the effectiveness of MGAug in image analysis tasks (IAT) by developing a joint learning framework. It has been shown in [65, 6, 7] that these two tasks can be mutually beneficial. The MGAug effectively synthesizes image samples to augment the training dataset, while the IAT can provide extra information on rectifying the learned distributions of latent geometric spaces. An overview of our proposed network architecture is shown in Fig. 2.
Figure 2: An overview of our model MGAug for image analysis tasks.
Network design of MGAug.
MGAug is mainly designed based on the previously introduced inference model, where the approximate posteriors are simulated by neural networks. Analogous to [32], we let the encoder of MGAug, parameterized by $\beta$, approximate the posterior as a mixture of multivariate Gaussians. In the latent space $\mathbf{z}$, we use non-linear hidden layers to learn the parameters (i.e., means and covariances) of the mixture Gaussians. The decoder, parameterized by $\theta$, outputs augmented velocity fields $\mathbf{v}_{n}$, followed by generating their associated deformations $\phi_{n}$ (via Eq. (3)) and augmented deformed images. In this paper, we adopt UNet architecture [50] as the network backbone for MGAug. However, other networks such as UNet++ [66] and TransUNet [8] can be easily applied.
Defining $\Theta=(\mathbf{J},\beta,\theta,\mathbf{x},\mathbf{z},\mathbf{v},\boldsymbol{\epsilon},I)$ as posterior parameters, we formulate the network loss of MGAug as a combination of dissimilarity between deformed images and templates, a regularity to ensure the smoothness of augmenting transformations, and the variational inference objective as
@@ III-C Joint Learning of MAGug With Image Analysis
$+\frac{1}{2}(\mathcal{L}\mathbf{v}(\beta,\theta),\mathbf{v}(\beta,\theta))+\text{reg}(\beta,\theta),$
where $\text{reg}(\cdot)$ denotes network parameter regularizations.
Network design of IAT.
@@
The total loss of our joint learning framework of MGAug and IAT is $\ell=\ell_{\text{MGAug}}+\ell_{\text{IAT}}$. We employ an alternative optimization scheme [46, 6] to minimize the loss by alternating between the training of the MGAug network and the IAT network without sharing weights, making it an end-to-end network architecture [65, 7].

### v3

@@ Abstract
Geometric transformations have been widely used to augment the size of training images. Existing methods often assume a unimodal distribution of the underlying transformations between images, which limits their power when data with multimodal distributions occur. In this paper, we propose a novel model, Multimodal Geometric Augmentation (MGAug), that for the first time generates augmenting transformations in a multimodal latent space of geometric deformations. To achieve this, we first develop a deep network that embeds the learning of latent geometric spaces of diffeomorphic transformations (a.k.a. diffeomorphisms) in a variational autoencoder (VAE). A mixture of multivariate Gaussians is formulated in the tangent space of diffeomorphisms and serves as a prior to approximate the hidden distribution of image transformations. We then augment the original training dataset by deforming images using randomly sampled transformations from the learned multimodal latent space of VAE. To validate the efficiency of our model, we jointly learn the augmentation strategy with two distinct domain-specific tasks: multi-class classification on 2D synthetic datasets and segmentation on real 3D brain magnetic resonance images (MRIs). We also compare MGAug with state-of-the-art transformation-based image augmentation algorithms. Experimental results show that our proposed approach outperforms all baselines by significantly improved prediction accuracy. Our code is publicly available at GitHub.
@@ I Introduction
Recent advances in deep learning (DL) have led to remarkable progress in image analysis tasks, including but not limited to image classification [38, 64, 27], segmentation [28, 10], object detection [35, 52], and 3D reconstruction [32]. The literature has shown that deep neural networks (DNNs) are most powerful when trained on vast amounts of manually labeled data. Unfortunately, many image applications have relatively small annotated datasets available, due to privacy concerns, regulations, costs, and incidents of disease. The DNNs, in such cases, tend to overfit due to their poor generalizations to large data variances. Data augmentation (DA) is a simple, yet efficient and commonly used strategy to overcome this issue by synthesizing more training examples to artificially increase the dataset [11, 72, 75, 58]. It is often employed as a preprocessing step applied to the training data to expose the DNNs to various types of approximated invariance; hence reducing the problem of overfitting. In the image domain, common augmentation strategies include adding variations to image intensities [7, 8, 19, 26, 73], or deforming the image by randomly generated geometric transformations (i.e., translation, rotation, or scaling) [47]. While intensity-based augmentation methods are particularly valuable for tasks that rely heavily on color and texture information, geometric transformations generate diverse, realistic scenarios introducing structural variations in the training data. This diversity in training data leads to improved generalization, as the model learns to recognize key features regardless of their spatial arrangement. In this paper, we focus on DA, particularly in geometric spaces that characterize deformable shape (changes) of objects presented in images [34, 4, 69, 67, 13].
@@
In practice, geometric DA is usually a manual process where a human specifies a set of affine or deformable transformations believed to make DL models invariant [24, 66]. However, such methods can pose a high risk of producing unrealistic or implausible variations of the image. For example, rotating an image 180 degrees makes sense for natural images, but produces ambiguities on digital numbers (i.e., the digits ‘6’ and ‘9’). In medical domain, randomly generated deformations may transform images from a healthy to a diseased group. To alleviate this issue, researchers have developed a group of automatic DA methods that learn an optimal augmenting policy from the data itself [11, 42, 75, 1, 49].
Despite showing encouraging results on relatively easy datasets (e.g., $28^{2}$ MNIST), they do not appear to be applicable to more complex datasets with large structural and geometric variances. Later, generative models (e.g., VAE) have been extensively used as a practical and effective solution to capture complex geometric variances for the purpose of DA in low-data regimes [46, 41, 53, 61]. The rationale for using VAEs in DA lies in their ability to learn and represent latent features of the data. By sampling from the learned latent space, VAEs can generate diverse and realistic synthetic data samples beyond the training dataset. Besides, in the context of DA, VAEs are not typically tasked with learning the entire data distribution but rather with generating synthetic data samples consistent with the original dataset’s distribution [5, 6, 49].
Another common approach is to learn a probabilistic distribution of a set of deformable transformations derived from the training images [22, 37, 14, 72, 51, 58]. However, these methods assume a unimodal distribution of the underlying transformations between images. They do not provide an appropriate way to learn the full space of geometric transformations among the training data. This greatly limits their power when multimodal data that naturally has a distribution of multiple modes, for example, time-series data/longitudinal images [15, 39], or images with sub-populations (i.e., population studies that involve healthy vs. disease groups [31, 39]).
@@
In this paper, we present a novel model, MGAug, that for the first time learns augmenting transformations in a multimodal latent space of geometric deformations. Inspired by works that have utilized mixture models to capture the underlying distribution of deformations from groupwise images, leading to improved analysis for population studies [68, 71, 43, 25], our MGAug features the main advantages of
@@
To achieve our goal, we first develop a deep learning network that embeds the learning of latent geometric spaces of diffeomorphic transformations via template-based image registration in a VAE. A mixture of multivariate Gaussians will be defined in the tangent space of diffeomorphisms and serve as a prior in the latent space of VAE. We then augment the original training dataset by deforming images using decoded velocity fields that are randomly sampled from the learned multimodal distributions. We demonstrate the effectiveness of our proposed model in two distinct domain-specific tasks: multi-class classification and segmentation on both 2D synthetic data and real 3D brain magnetic resonance images (MRIs), respectively. We also compare MGAug with the state-of-the-art transformation-based augmentation algorithms [51, 14]. Experimental results show that our method outperforms all baseline algorithms by significantly improved prediction accuracy.
The remainder of this paper is organized as follows: Sec. II discusses the background of geometric deformations via template-based image registration, Sec. III-A introduces the methodological development of our proposed augmentation algorithm MGAug along with joint optimization scheme, Sec. IV discusses the experimental evaluations that validate the effectiveness of our model on different image analysis tasks, and Sec. VI concludes with potential future directions.
@@ II Background: Geometric Deformations via Template-based Image Registration
$\quad\text{with}\quad\frac{d\phi_{n}(t)}{dt}=v_{n}(t)\circ\phi_{n}(t),$ (2)
@@
Numerous research works have defined a prior on the initial velocity field, $\tilde{v}_{n}$, as a multivariate Gaussian distribution to ensure the smoothness of the geodesic path [63, 69, 70]. This prior is typically formulated as
$p({v}_{n})=\frac{1}{(2\pi)^{\frac{M}{2}}|{L}^{-1}|^{\frac{1}{2}}}\exp{\left(-\frac{1}{2}({L}{v}_{n},{v}_{n})\right)},$ (5)
where $M$ denotes the total number of voxels of initial velocities, and ||$\cdot$|| ⋅ | is the matrix determinant. Note that here we consider the input images to be measured on a discretized grid. That is, images are elements of the finite-dimensional Euclidean space. We will also consider velocity fields and the resulting diffeomorphisms to be defined on the discrete grid. This allows us to work with a well-defined probability density function while maintaining the smoothness properties of the velocity fields through the use of the differential operator $L$. In practice, we employ a discretized operator $L$ (as previously defined in Eq. (2)) [4, 62, 69]. The $|L^{-1}|$ refers to the determinant of the inverse of the discretized differential operator.
@@ III Our Method: MGAug
This section presents a novel algorithm, MGAug, that learns a multimodal latent distribution of geometric transformations represented by the initial velocity fields $v_{0}$. We first introduce a generative model in latent geometric spaces for DA based on VAE, followed by an approximated variational inference. The augmented data is then used to train various image analysis tasks, i.e., image classification or segmentation. Similar to [72, 9], we employ a joint learning scheme to optimize the MGAug and image analysis tasks (IAT) network parameters. To simplify the notation, we omit the time index of $v_{0}$ in the following sections.
@@ III-A Generative Model in Latent Multimodal Geometric Spaces
Given a group of $N$ input images $\mathbf{J}$ with a template $I$, we assume the latent geometric transformations $\mathbf{x}$ are generated from a number of $C$ clusters. The distribution of $\mathbf{x}$ can be formulated as a Gaussian mixture with means $\boldsymbol{\mu}$ and variances $\mathbf{\Sigma}$, specified by a VAE encoder parameterized by $\beta$, i.e., $p_{\theta}(\mathbf{x}|{\boldsymbol{\epsilon}})$, where $\boldsymbol{\epsilon}$ is a random number generated from $\mathcal{N}(0,\mathbb{I})$. That is to say, the VAE encoder outputs a set of $C$ means $\boldsymbol{\mu}$ and $C$ variances $\mathbf{\Sigma}$. A one-hot vector $\mathbf{z}$ is sampled from the mixing probability $\mathbf{\boldsymbol{\pi}}\in[0,1]$ with $\sum_{c=1}^{C}\mathbf{\boldsymbol{\pi}}^{c}=1$, which chooses one component from the Gaussian mixture. In this paper, we set the parameter $\mathbf{\boldsymbol{\pi}}=C^{-1}$ to make the variable $\mathbf{z}$ uniformly distributed, which is similar to [17]. The initial velocity field $\mathbf{v}$ is generated from a VAE decoder parameterized by $\theta$ and the continuous latent variable $\mathbf{x}$. The entire process to generate the observed image data $\mathbf{J}$ from a set of latent variables $\{\mathbf{v},\mathbf{x},\mathbf{\boldsymbol{\epsilon}},\mathbf{z}\}$ is:
@@
Similar to Eq. (5), we define the covariance matrix $\boldsymbol{\Sigma}$ as an inverse operator of the discretized differential operator, $L$, which is typically used in the registration regularity in Eq. (2) to guarantee the smoothness of transformations and preserve the topological structures of images [4].
@@ III-B Variational Lower Bound
We will train the generative model in Eq. (13) with the variational inference objective, i.e. the log-evidence lower bound (ELBO) [36]. With an assumption that the mean-field variational family $q(\mathbf{v},\mathbf{x},\mathbf{\boldsymbol{\epsilon}},\mathbf{z}\,|\,\mathbf{J})$ is an approximate to the posterior, we are able to formulate the ELBO as
@@
We factorize the mean-field variational family as
$\mathbf{q}=\prod_{i=1}^{n}q(\mathbf{v}_{i}|\mathbf{J}_{i},I,\boldsymbol{\lambda})\,q_{\psi_{\mathbf{x}}}(\mathbf{x}_{i}|\mathbf{v}_{i},\mathbf{J}_{i})\,q_{\psi_{\mathbf{\boldsymbol{\epsilon}}}}(\mathbf{\boldsymbol{\epsilon}}_{i}|\mathbf{v}_{i})\,p_{\beta}(\mathbf{z}_{i}|\mathbf{x}_{i},\mathbf{\boldsymbol{\epsilon}}_{i}),$ (15)
where $i$ denotes the $i$-th image. Each variational factor is parameterized with the recognition networks $\psi_{\mathbf{x}}$ and $\psi_{\mathbf{\boldsymbol{\epsilon}}}$ that output the parameters of the variational Gaussian distributions. Note that we will drop the index $i$ in the following sections for the purpose of notation simplicity.
Figure 2: An overview of our model MGAug for image analysis tasks.
@@
$-\mathbb{E}_{q(\mathbf{x}|\mathbf{J})\,q(\boldsymbol{\epsilon}|\mathbf{J})}[KL(p_{\beta}(\mathbf{z}|\mathbf{x},\boldsymbol{\epsilon})\,||\,p(\mathbf{z}))]$ (16)
III-C Joint Learning of MGAug with Image Analysis
We demonstrate the effectiveness of MGAug in image analysis tasks (IAT) by developing a joint learning framework. It has been shown in [72, 9] that these two tasks can be mutually beneficial. The MGAug effectively synthesizes image samples to augment the training dataset, while the IAT can provide extra information on rectifying the learned distributions of latent geometric spaces. An overview of our proposed network architecture is shown in Fig. 2.
III-C1 Network design of MGAug
Our MGAug is mainly designed based on the previously introduced inference model, where the approximate posteriors are simulated by neural networks. Analogous to [36], we let the encoder of MGAug, parameterized by $\beta$, approximate the posterior as a mixture of multivariate Gaussians. In the latent space $\mathbf{x}$, we use non-linear hidden layers to learn the parameters (i.e., means and covariances) of the mixture Gaussians. The decoder, parameterized by $\theta$, outputs augmented velocity fields $\mathbf{v}_{n}$, followed by generating their associated deformations $\phi_{n}$ (via Eq. (3)) and augmented deformed images. In this paper, we adopt UNet architecture [55] as the network backbone for MGAug. However, other networks such as UNet++ [74] and TransUNet [10] can be easily applied.
Defining $\Theta=(\mathbf{J},\beta,\theta,\mathbf{x},\mathbf{z},\mathbf{v},\boldsymbol{\epsilon},I)$ as posterior parameters, we formulate the network loss of MGAug as a combination of dissimilarity between deformed templates and input images, a regularity to ensure the smoothness of augmenting transformations, and the variational inference objective as
@@ III-C1 Network design of MGAug
$+\frac{1}{2}(L\mathbf{v}(\beta,\theta),\mathbf{v}(\beta,\theta))+\text{reg}(\beta,\theta),$
where $\text{reg}(\cdot)$ denotes network parameter regularization.
III-C2 Network design of IAT
@@
The total loss of our joint learning framework of MGAug and IAT is $\ell=\ell_{\text{MGAug}}+\ell_{\text{IAT}}$. We employ an alternative optimization scheme [50] to minimize the loss by alternating between the training of the MGAug network and the IAT network without sharing weights, making it an end-to-end network architecture [72, 9].
Input : A group of $N$ input images with an image template $I$, class labels $\hat{y}_{\mathbf{J}}$ or ground truth segmentation labels $\hat{Y}_{\mathbf{J}}$ and a number of iteration $r$.
Output : Predicted classification labels $y_{\mathbf{J}}$ or segmentations $Y_{\mathbf{J}}$.
1 repeat
2       for i = 1 to $r$ do
             /* Train MGAug */
3             Minimize the augmentation loss in Eq. (III-C1)
4            Generate augmented images (and propagate associated deformations to segmentation labels, for segmentation);
            /* Train classifier or segmentation network */
5
6            Mix augmented images and labels with ground truth labels, $\hat{y}_{\mathbf{J}}$ or $\hat{Y}_{\mathbf{J}}$, in training;
7            Minimize the classification loss in Eq. (17) or segmentation loss in Eq. (18);
8            Output the predicted class labels $y_{\mathbf{J}}$ or segmentations $Y_{\mathbf{J}}$;
9       end for
10
11until convergence, $|\Delta\ell|<\epsilon$;
Algorithm 1 Joint Training of MGAug with IAT.
A summary of our joint training of MGAug with IAT is presented in Alg. 1.